# Towards Next-Generation Recommender Systems: A Benchmark for Personalized Recommendation Assistant with LLMs

### v1

@@ 3.1.2. Evaluation Metrics
  • Fail to Recommend (FTR): This metric measures the proportion of queries where the model fails to generate recommended items. A lower FTR is preferred, indicating the model can generate a list of items. However, in Misinformed Condition-based Query, a higher FTR shows the model’s ability to identify and reject incorrect information, thereby avoiding wrong recommendations.
@@ 3.2. Results on Condition-based Query
Observation 1. For different LLMs, GPT-4o and DeepSeek-R1 outperform other models. As shown in Table 2, GPT-4o achieves the highest Precision and the second-highest CMR average score on both the movie and book datasets, while DeepSeek-R1 ranks first in Recall and second in Precision. Notably, models with advanced reasoning capabilities, such as DeepSeek-R1, excel in handling query types that require reasoning. For example, in Implicit Condition Query, DeepSeek-R1 exhibits a significantly smaller drop in performance compared to other models when transitioning from explicit to implicit conditions. This positions it as the leader in Implicit Condition Query performance. It suggests that the advanced reasoning ability enables the LLM-based assistant to infer implicit attributes not explicitly mentioned in the conditions, leading to more accurate recommendations. There is a related case study in Appendix D.2.
@@
Figure 4. Performance on Condition-based Query with different number of conditions.
@@
Observation 4. For the effect of user-item interaction history, incorporating them enhances recommendation quality in two key ways. The effect of incorporating user-item interaction history is shown in Figure 5. First, incorporating user-item history improves Precision across all query types by generating a more personalized recommendation list. The reason can be that queries with limited conditions (e.g., “movies directed by Nolan”) often yield a large candidate pool. By leveraging historical interactions, the model can prioritize select items that align with the user’s preferences and satisfy the query conditions, effectively filtering irrelevant candidates. We provide examples of both using and not using history in the Appendix D.3.
@@
Second, incorporating user history does not always improve CMR. By analyzing the output of the LLM assistant, we note that adding history can also introduce “distractor” items that reduce the model’s strict condition adherence. For example, the LLM assistant may recommend “The Prestige” for “Nolan films” due to the user’s fantasy genre bias in history. This suggests that incorporating history may divert the model’s attention, making it more challenging to strictly adhere to the given conditions. These results highlight an important trade-off: while incorporating history can enhance the model’s understanding of user preferences, it may also hinder the model’s ability to closely align the specific requirements in the query. Consequently, balancing the influence of historical data with the need for precise condition adherence is crucial for optimizing performance in recommendation assistants.
@@ 3.3. Results on User Profile-based Query
Observation 5. For different LLMs, Gemini-1.5 Pro and DeepSeek-R1 demonstrated better performance compared to other models. As shown in Table 3, larger and more advanced models, such as DeepSeek-R1 and Gemini-1.5-Pro, consistently achieve higher Precision and Recall. In contrast, smaller models like GPT-4o-mini tend to lag behind. From the FTR (Failure to Recommend) metrics perspective, most models exhibit a low failure rate, indicating their ability to effectively follow instructions and generate recommendation lists in the required format.
@@
Domain Model Interest-based Query Demographics-based Query Average
@@
Observation 6. For different query types, Demographics-based Query exhibit lower Recall than Interest-based Query but similar Precision. Across all models, Recall for Demographics-based Query is generally lower than for Interest-based Query. This suggests that LLMs struggle to identify relevant recommendations when relying on user demographics. The difference in Recall can be attributed to the distinct ways in which these queries are constructed. Interest-based Query leverages clear interest, making it easier for models to generate relevant recommendations. In contrast, Demographics-based Query requires recommendation assistants to infer preferences from broader demographic and behavioral trends, which introduces more variability and ambiguity.
@@
Observation 7. For Interest-based Query, queries constructed from less prevalent interests tend to be more challenging. As shown in Figure 6, we categorized the queries into five groups based on their popularity and conducted experiments on the movie and book datasets. It can be observed that interests that are shared by more users tend to yield higher Precision and Recall. This is likely because widely shared interests are more easily recognizable and interpretable by LLMs, leading to better performance. In contrast, the book domain shows an opposite trend. This is due to the nature of book data and the evaluation process. Books often have multiple editions or publishers, and popular titles tend to have more variants. Therefore, the popular book (e.g., “Dune”) can suffer lower metrics because LLMs confuse editions/publishers (e.g., “Dune: 1965” vs. “2021 adaptation”). For less popular books, fewer variants exist, making it easier for LLMs to provide exact matches, which leads to higher performance metrics.
Observation 8. For Demographics-based Query, the performances of LLMs show notable variations based on different user demographics. For instance, as figure 9 in the Appendix E shows, we note that models exhibit higher accuracy for female users. It reflects more consistent preference patterns of female users (e.g., stable genre preferences like romantic comedies), which allows the model to make accurate recommendations. Regarding occupation, the model performs best for users in sales/marketing roles, likely because these professions are associated with more consistent and recognizable patterns in user behavior. Age-wise, LLMs perform better regarding Recall for people aged 50–55. This is likely because their preferences are more focused and less affected by rapidly changing popular culture compared to younger users, who tend to have more diverse preferences. For users aged 56 and above, the model’s effectiveness is not as strong as it is for the 50–55 age group, possibly because older users engage less frequently in online activities, such as rating or reviewing movies. As a result, there is less data about this age group in the training set of LLMs, which makes it more challenging for models to understand their preferences.
@@ 4. Related Work
Recent advances in LLM-based recommender systems have led to the development of various benchmarks to evaluate their capabilities (Liu et al., 2023; Jiang et al., 2024; Tan et al., 2025; Xue et al., 2025; Zhang et al., 2023a). Early efforts like LLMRec (Liu et al., 2023) focus on traditional recommendation tasks (e.g., rating prediction, sequential recommendation), establishing baselines for LLMs against classical methods. In addition to traditional recommendation tasks, recent studies have increasingly focused on evaluating the unique characteristics of LLM for recommendations. For instance, one recent study (Jiang et al., 2024) introduces a multi-dimensional framework that assesses not only utility and novelty but also LLM’s specific behaviors like history length sensitivity (how recommendations perform with sparse user histories) and hallucination (generation of non-existent items). Similarly, PerRecBench (Tan et al., 2025) isolates personalization accuracy by eliminating biases in users and item ratings, testing whether LLMs infer user preferences.
Meanwhile, with the rise of LLMs, varying benchmarks have been proposed to evaluate their capabilities as assistants. For example, Gaia (Mialon et al., 2023) measures LLM’s capabilities as a general-purpose AI assistant. In addition, domain-specific frameworks like VoiceBench (for voice assistant) (Deng et al., 2024) and Mobile-Bench (for mobile assistant) (Chen et al., 2024) assess the ability of LLM as the specific assistant. However, despite progress in both recommendation and assistant evaluation, there remains a gap in assessing LLMs as personalized recommendation assistants. To bridge this gap, we introduce RecBench+, a novel benchmark designed to evaluate LLMs’ potential as personalized recommendation assistants. Due to the space limitation, some studies about RecSys studies are reviewed in Appendix A.
@@ 5. Conclusion
In this paper, we introduce a new paradigm for next-generation recommender systems and propose a novel benchmark dataset, RecBench+, to evaluate the potential of LLMs as personalized recommendation assistants. Our benchmark includes five specially designed query types that comprehensively assess LLM-based assistants’ capabilities in handling complex user queries in real-world scenarios. We conduct extensive experiments on a diverse set of LLMs, including those with advanced reasoning capabilities. Through these experiments, we present eight detailed observations that shed light on the strengths and limitations of LLMs in serving as personalized recommendation assistants. We hope our benchmark and findings lay a solid foundation for future research and development in this emerging domain. We provide more discussion about future directions in Appendix B.
Appendix A More Related Work
RecSys as intelligent filtering tools have gained rapid development in the past decades. Early methods such as NCF (He et al., 2017) and SVD++ (Koren, 2008) leverage collaborative filtering (CF) information to generate recommendations. With the evolution of deep neural networks, Graph Neural Networks (GNNs) have dominated RecSys due to their ability to model high-order users’ CF information (He et al., 2020; Fan et al., 2019; Ying et al., 2018). Typical studies like LightGCN (He et al., 2020) and GraphRec (Fan et al., 2019) introduce GNN to aggregate neighborhood information for user and item embeddings to improve recommendations.
Recently, with the emergence of LLMs, various LLM-based RecSys have attracted widespread attention. With powerful reasoning and language understanding capabilities, LLMs can delve deeper into the semantic understanding of user preferences and item characteristics (Geng et al., 2022; Bao et al., 2023; Zhang et al., 2023b; Qu et al., 2024; Wang et al., 2025). For example, P5 (Geng et al., 2022) unified various recommendation tasks as the text generation task. TALLRec (Bao et al., 2023) and CoLLM (Zhang et al., 2023b) further tuned on user-item interaction data to compensate for inadequate recommendation data during the pre-training phase. Although existing recommendation frameworks can address conventional tasks like rating prediction (Geng et al., 2022; Koren, 2008) and sequential recommendation (Bao et al., 2023; Zhang et al., 2023b; Wang et al., 2025), their ability to handle diverse, complex, and multi-faceted user needs remains an open question. Most of them directly transform the datasets from traditional recommendation paradigm to the interactive paradigm in the era of LLMs, which is inappropriate for practical scenarios and insufficient for evaluating recommendation performance. Note that modern users increasingly require personalized and context-aware recommendations that deeply understand multifaceted user needs. To fill this research gap, we propose a holistic benchmark to systematically evaluate the potential of various advanced LLMs to serve as recommendation assistants.
Appendix B Future Directions
In this study, we proposed a novel paradigm for evaluating the LLM-based recommendation. We systematically examined the potential of widely used LLM backbones to serve as personalized recommendation assistants in our RecBench+ benchmark dataset. Our findings showed that LLMs already exhibit preliminary capabilities as personalized recommendation assistants, yet their full potential remains under-explored. This highlights the need for further investigation to fully unlock the power of LLMs in personalized recommendation tasks. Several future directions present exciting opportunities to enhance and expand the capabilities of LLM-based recommendation assistants.
One important direction to explore is the effectiveness of LLMs in real-world recommendation scenarios, especially after specialized fine-tuning. Our study shows that pure LLM backbones possess initial capabilities as personalized recommendation assistants, and different LLMs exhibit varying strengths in handling specific types of user queries. For instance, models like GPT-4o and DeepSeek-R1 demonstrate strong performance in handling user queries with explicit conditions for expected items, whereas Gemini-1.5-Pro excels in processing queries without explicit conditions, requiring user profile understanding. Therefore, it is necessary to explore how specialized fine-tuning can further optimize their performance as personalized recommendation assistants. This includes evaluating whether such training can lead to improvements across different tasks, the extent of these improvements, and whether certain models benefit more than others from specialized training.
Another key direction for improvement is integrating external tools to enhance the performance of LLM-based recommendation assistants. External tools, such as search engines and knowledge bases, can provide valuable up-to-date information and real-world context that LLMs may lack due to the static nature of their training data. By incorporating these tools, recommendation assistants can better understand the world and current events, enabling them to provide more timely, intelligent, and customized recommendations. This integration could greatly enhance the alignment of recommendations with users’ evolving needs and preferences, delivering more personalized and intelligent recommendations.
Finally, in domains where LLMs lack specialized knowledge—such as in e-commerce recommendations—combining assistants with techniques like Retrieval-Augmented Generation (RAG) or other hybrid models could prove highly effective. LLMs typically do not have access to domain-specific knowledge, such as detailed item pools or up-to-date product information, making them ill-equipped to provide accurate recommendations in these areas. RAG techniques may address this limitation by using external retrieval mechanisms to supplement the generative capabilities of LLMs. This allows the model to dynamically query databases or corpora of item information, enhancing the accuracy and contextual relevance of the recommendations. Without such integration, LLMs alone may struggle to offer satisfactory recommendations in specialized domains.
Appendix C Query Example
Table 4. Examples of queries from different types in our RecBench+ benchmark.
Task Query
Explicit Condition Query I’m really interested in classic films and would love to watch something that showcases Charlie Chaplin’s legendary comedic talent. Additionally, I’ve heard that Roland Totheroh’s cinematography adds an exceptional visual quality to movies. If you could point me in the direction of films that include both of these elements, I’d greatly appreciate it!
Implicit Condition Query I recently watched Clockers (1995) and Bamboozled (2000), and I was really impressed by the direction in both films. I’m eager to explore more works from the director, as I found their storytelling style and vision very engaging. If you could suggest other films associated with this director, that would be fantastic.
Misinformed Condition Query I recently watched Lorenzo’s Oil and was really impressed by the cinematography done by Mac Ahlberg. I’m interested in finding more films that showcase his cinematographic style. I also remember seeing his work in Beyond Rangoon, so if there are any other movies he contributed to, I’d love to check them out!
Interest-based Query I’m fond of romantic and dramatic films from the golden age of Hollywood like ‘Roman Holiday’ and ‘My Fair Lady’. Are there any other dramatic romances from that period you would recommend?
Demopraphics-based Query I’m a psychology professor and I’m looking for movies that delve into human emotions and relationships. Have you got any?
Examples of different categories of queries in our benchmark are shown in Table 4.
Appendix D Case Study
We carefully examined LLM-based recommendation assistants’ outputs and identified three key factors significantly influencing the recommendation performance. In the following sections, we provide examples of each of these factors.
D.1. The Impact of LLM Knowledge on Performance
Model knowledge plays a crucial role in the quality and relevance of recommendations. In our evaluation, we observed that recommendation assistants often fail to generate suitable recommendations when they lack sufficient domain-specific knowledge, even if the query explicitly states conditions for the user’s expected items.
In the case presented, GPT-4o-mini fails to recommend movies featuring cinematography by Ellery Ryan, responding that it lacks the necessary information. In contrast, GPT-4o successfully recommends relevant films, such as I Love You Too (2010) and The Rage in Placid Lake (2003). This performance difference suggests that GPT-4o may possess a more extensive knowledge base or improved knowledge-recalling mechanisms, allowing it to generate appropriate recommendations where the other fails.
D.1 Case for Impact of LLM Knowledge Query: I’m looking for a movie with stunning visuals, specifically those that have the unique touch of cinematography by Ellery Ryan. I’ve heard that his work enhances storytelling, and I’d love to experience it firsthand. Output (GPT-4o-mini): I’m sorry, but I am unable to provide movie recommendations specifically featuring cinematography by Ellery Ryan, as I do not have data on his work. Output (GPT-4o): I Love You Too (2010), The Rage in Placid Lake (2003)
D.2. The Impact of Advanced Reasoning Capability on Performance
Advanced reasoning is critical in improving the robustness and accuracy of recommendation assistants, especially when a user’s query contains factual errors that could serve as a distraction. LLMs with strong reasoning capabilities are more likely to verify information, resolve inconsistencies, and generate appropriate recommendations.
In the first case study, the user inquires about movies featuring cinematography by John Blick, referencing The Mirror Has Two Faces (1996), which is incorrect. Recommendation assistants lacking a clear reasoning process would attempt to match John Blick with known cinematographers, leading to incorrect recommendations. However, DeepSeek-R1, which possesses advanced reasoning capability, successfully identifies that John Blick was not the cinematographer of the mentioned film, instead recognizing the correct individual, Dante Spinotti. By reasoning through the discrepancy, the model does not give any movies directed by John Blick.
D.2 Case 1 for Impact of Advanced Reasoning Capability Query: I recently watched The Mirror Has Two Faces and was really impressed by John Blick’s cinematography. I’m interested in finding more films that feature his work behind the camera. If you know of any other movies that he did the cinematography for, I’d love to check them out! Ground Truth: Falling Down (1993), Speed (1994), The Devil’s Advocate (1997), U.S. Marshals (1998); or ”None” Output (DeepSeek-V3): The Last Seduction, One False Move, The Underneath, Albino Alligator, The Lookout Output (DeepSeek-R1): None Reasoning (DeepSeek-R1): … first, I need to confirm if John Blick is the actual cinematographer for that movie. Let me check my database. Ah, T̈he Mirror Has Two Faces(̈1996) was actually cinematographed by Dante Spinotti, not John Blick. So there’s a mistake here. Maybe the user misremembered the name. Alternatively, maybe John Blick is a lesser-known DP or a pseudonym? Let me cross-reference. No, seems no record of John Blick working on that film…Well, given the instructions, I need to answer with the movies, no explanations. So since there are none, the field is empty.
D.2 Case 2 for Impact of Advanced Reasoning Capability Query: I recently enjoyed the cinematography in Taps (1981) and Absence of Malice (1981), both of which showcased Scott Ambrozy’s incredible work. I’m eager to find more films that feature his cinematography style. Any recommendations? Ground Truth: 3 Days of the Condor (a.k.a. Three Days of the Condor) (1975), Network (1976), The Exorcist (1973), The Stepford Wives (1975) ; or ”None” Output (DeepSeek-V3): The Chosen (1981), Four Friends (1981), Split Image (1982), Heart Like a Wheel (1983) Output (DeepSeek-R1): The French Connection (1971), The Exorcist (1973), Network (1976), Three Days of the Condor (1975), The Electric Horseman (1979), Straight Time (1978), The Prince of the City (1981), Wyatt Earp (1994). Note: The films listed above feature the cinematography of Owen Roizman, who shot Taps and Absence of Malice. Scott Ambrozy appears to be a confusion or typo in the query. Reasoning (DeepSeek-R1): … Wait, first, who is Scott Ambrozy? I know some big cinematographers, but this name isn’t ringing a bell. Maybe I should verify if there’s a Scott Ambrozy in the industry. Let me think. Maybe the user misspelled the name? Looking up Taps on IMDb, yeah, Owen Roizman is the cinematographer for Taps. Absence of Malice also lists Owen Roizman. So Scott Ambrozy might be a confusion with another name. So the user probably meant Owen Roizman. So to answer correctly, I should recommend other films by Owen Roizman. Since the user wants movies with his cinematography style, I should list those. Roizman’s notable works include…
A similar pattern emerges in the second case, where the user asks for films by Scott Ambrozy, citing Taps (1981) and Absence of Malice (1981). The user is actually seeking recommendations for movies with the same cinematographer as these two films. However, the model DeepSeek-V3, lacking strong reasoning capabilities, tries to list movies by Scott Ambrozy, resulting in incorrect recommendations. In contrast, DeepSeek-R1 takes a different approach. It questions whether Scott Ambrozy is actually a cinematographer first. Subsequently, it identifies that the actual cinematographer of these films is Owen Roizman. Finally, DeepSeek-R1 infers that the user probably meant to ask about Roizman’s work, providing correct recommendations.
These examples highlight that the ability to evaluate and refine its reasoning process significantly enhances recommendation accuracy. Instead of passively following instructions, models with advanced reasoning capabilities proactively identify inconsistencies, validate claims, and generate factually sound responses. This improvement is particularly impactful in real-world applications, where user inputs often contain typos, misinformation, or ambiguous phrasing.
D.3. The Impact of User-Item Interaction History on Performance
User-item interaction history is crucial for personalized recommendation assistants to align their recommendation outputs with individual user preferences. Assistants lacking access to historical data may generate recommendations that meet the user’s criteria yet fail to fully align with their preferences. Our case study demonstrates how historical preferences influence recommendation accuracy.
D.3 Case for Impact of User-Item Interaction History Query: I’m looking for a nostalgic film experience, ideally featuring the talented duo of Karen Dotrice and Matthew Garber. Their performances always bring a sense of charm and whimsy that I adore, so I’d appreciate any recommendations that showcase them together! User History: Starship Troopers (1997), You’ve Got Mail (1998), Beetlejuice (1988), Pocahontas (1995), Beauty and the Beast (1991), Rescuers, The (1977), On Golden Pond (1981), Star Wars: Episode V - The Empire Strikes Back (1980), Bowfinger (1999) Ground Truth: Mary Poppins (1964), The Gnome-Mobile (1967) Output (GPT-4o, w/o history): The Three Lives of Thomasina Output (GPT-4o, with history): Mary Poppins (1964), The Gnome-Mobile (1967) Note: *Mary Poppins* (1964) is a strong recommendation as it blends Disney’s charm with whimsical fantasy, aligning with the user’s preference for animated classics like *Beauty and the Beast* and *The Rescuers*. Its mix of adventure and magic also echoes the fantastical elements of *Beetlejuice* and the nostalgia of *You’ve Got Mail*. *The Gnome-Mobile* (1967) is another fitting choice, offering a lighthearted adventure with Karen Dotrice and Matthew Garber, while its humor parallels *Bowfinger*. In contrast, *The Three Lives of Thomasina* (1963) leans more toward sentimentality and drama, deviating from user’s favored whimsical and adventurous films.
When the LLM (GPT-4o) doesn’t have access to the user’s history, it recommends The Three Lives of Thomasina (1963). This movie features Karen Dotrice, but it’s more about sentimentality and drama. The recommendation misses the user’s preference for whimsical, adventurous, and nostalgic films. However, when user history is incorporated, the LLM-based assistant correctly recommends Mary Poppins (1964) and v. These movies match the user’s requirements for Karen Dotrice and Matthew Garber’s performances and fit the themes and styles of the films the user has enjoyed before.
This performance gap underscores the importance of incorporating past interactions into recommendation assistants. Without access to history, assistants mainly rely on the conditions for items in the query to provide recommendations, which can result in recommendations that meet the basic criteria but don’t fully satisfy the user’s preferences. In contrast, integrating user history allows personalized recommendation assistants to infer underlying preferences and provide recommendations that are not only conditionally relevant but also more aligned with user preferences. By analyzing past interactions, the assistant can refine its understanding of user preferences, leading to more personalized and contextually relevant suggestions. This highlights the importance of leveraging historical data to enhance recommendation accuracy and user satisfaction.
Table 5. Results on Condition-based Query with K=5
Model Explicit Condition Implicit Condition Misinformed Condition Average
Precision Recall CMR FTR Precision Recall CMR FTR Precision Recall CMR FTR Precision Recall CMR
GPT-4o 0.233 0.436 0.657 0.002 0.123 0.245 0.266 0.003 0.024 0.052 0.077 0.100 0.127 0.244 0.333
Gemini 0.223 0.416 0.607 0.002 0.099 0.205 0.195 0.003 0.023 0.048 0.062 0.003 0.115 0.223 0.288
Claude 0.234 0.436 0.647 0.002 0.126 0.247 0.277 0.010 0.035 0.075 0.105 0.063 0.132 0.253 0.343
Figure 7. Performances on Condition-based queries with different number of conditions when K=5.
Figure 8. The effect of incorporating user-item history when K=5.
Appendix E Visualization of Performance of Queries Constructed based on Different User Demographics
For Demographics-based Queries, the performances of LLMs show notable variations based on different user demographics. We visualize the trends in figure 9.
Figure 9. Average Recall of queries constructed based on different user demographics.
Appendix F Experiments with Fixed K for Recommendation Performance
We evaluated the performance of several well-performing models by testing their ability to generate top K recommended items. The prompt used for this experiment involved modifying the one in  G from “Please output the recommended movies directly” to “Please output {K} recommended movies directly.” Table  5 presents the experimental results on the movie dataset. The findings are consistent with the observations in the main text: the LLMs perform best on Explicit Condition Query, and their performance gradually decreases on Implicit Condition and Misinformed Condition Query.
Furthermore, we analyzed the relationship between model performance, the number of conditions, and user history. The results are illustrated in Figures 7 and 8, and the observed trends align with the observations in the main text.
Appendix G Prompts for Model Testing

### v2

@@ 2.1.2. Evaluation Metrics
  • Fail to Recommend (FTR): This metric measures the proportion of queries where the model fails to generate recommended items. A lower FTR is preferred, indicating the model can generate a list of items. However, in Misinformed Condition-based Query, a higher FTR shows the model’s ability to identify incorrect information, thereby avoiding wrong recommendations.
@@ 2.2. Results on Condition-based Query
Observation 1. For different LLMs, GPT-4o and DeepSeek-R1 outperform other models. As shown in Table 2, GPT-4o achieves the highest Precision and the second-highest CMR average score on both the movie and book datasets, while DeepSeek-R1 ranks first in Recall and second in Precision. Notably, models with advanced reasoning capabilities, such as DeepSeek-R1, excel in handling query types that require reasoning. For example, in Implicit Condition Query, DeepSeek-R1 exhibits a significantly smaller drop in performance compared to other models when transitioning from explicit to implicit conditions. This positions it as the leader in Implicit Condition Query performance. It suggests that the advanced reasoning ability enables the LLM-based assistant to infer implicit attributes not explicitly mentioned in the conditions, leading to more accurate recommendations. There is a case study in Appendix B.
@@
Figure 4. The effect of number of conditions.
Observation 4.
@@
For the effect of user-item interaction history, incorporating them enhances recommendation quality in two key ways. The effect of incorporating user-item interaction history is shown in Figure 5. First, incorporating user-item history improves Precision across all query types by generating a more personalized recommendation list. The reason can be that queries with limited conditions (e.g., “movies directed by Nolan”) often yield a large candidate pool. By leveraging historical interactions, the model can prioritize select items that align with the user’s preferences and satisfy the query conditions, effectively filtering irrelevant candidates.
Second, incorporating user history does not always improve CMR. By analyzing the output of the LLM assistant, we note that adding history can also introduce “distractor” items that reduce the model’s strict condition adherence. For example, the LLM assistant may recommend “The Prestige” for “Nolan films” due to the user’s fantasy genre bias in history. This suggests that incorporating history may divert the model’s attention, making it more challenging to strictly adhere to the given conditions. These results highlight an important trade-off: while incorporating history can enhance the model’s understanding of user preferences, it may also hinder the model’s ability to closely align with the specific requirements in the query. Consequently, balancing the influence of historical data with the need for precise condition adherence is crucial for optimizing performance in recommendation assistants.
Table 3. Results on Condition-based Query with K=5
Model Explicit Implicit Misinformed Average
Recall CMR Recall CMR Recall CMR Recall CMR
GPT-4o 0.436 0.657 0.245 0.266 0.052 0.077 0.244 0.333
Gemini 0.416 0.607 0.205 0.195 0.048 0.062 0.223 0.288
Claude 0.436 0.647 0.247 0.277 0.075 0.105 0.253 0.343
Observation 5. Requiring the LLM to generate K recommendations aligns with previous observations. For Condition-based Query, there are samples where the number of items that satisfy the conditions in the query may be fewer than K. In such cases, requesting the model to output K recommendations could lead to suboptimal results. Therefore, in the main experiment presented in Table 2, we did not require the model to generate K recommendations. However, to further explore this aspect, we also conducted additional experiments in which the LLM was instructed to output K items. The results, shown in Table 3, are based on the movie dataset and align with previous findings: LLMs perform best on Explicit Condition Queries, with performance gradually declining for Implicit Condition and Misinformed Condition Queries.
Table 4. Fine-tuning performance on Condition-based Query
Domain Model Fine-tuning Explicit Implicit Misinformed
Movie Qwen-2.5-3B - 0.019 0.002 0.004
SFT 0.391 0.255 0.296
RFT 0.178 0.119 0.089
SFT+RFT 0.410 0.262 0.307
Llama-3.2-3B - 0.018 0.003 0.003
SFT 0.285 0.176 0.186
RFT 0.191 0.124 0.080
SFT+RFT 0.293 0.181 0.195
Book Qwen-2.5-3B - 0.012 0.007 0.002
SFT 0.136 0.251 0.249
RFT 0.055 0.090 0.081
SFT+RFT 0.139 0.262 0.264
Llama-3.2-3B - 0.006 0.004 0.002
SFT 0.088 0.221 0.181
RFT 0.056 0.092 0.090
SFT+RFT 0.112 0.237 0.190
Observation 6. For the performance of fine-tuning, both SFT and RFT significantly enhance the model’s recommendation performance. Given the recent interest in using RFT to improve model reasoning capabilities (Shao et al., 2024; Xu et al., 2025), we aimed to explore its potential in this context. We applied full-parameter SFT and RFT (based on the GRPO (Shao et al., 2024) algorithm) to fine-tune the Qwen-2.5-3B-Instruct and LLaMA-3.2-3B-Instruct models, and evaluated recall performance across different query types. As shown in Table 4, both fine-tuning paradigms notably improved the models’ recommendation performances. However, models trained solely with RFT performed worse than those trained with SFT. This performance gap may be attributed to the lack of a warm-up phase for reasoning, which is essential for effective exploration during reinforcement learning (Cai et al., 2025). To validate this, we further divided the training process into two stages: first applying SFT as a warm-up, followed by RFT training. Experiment results demonstrate that this two-stage approach further improved LLMs’ performance. For example, in the movie’s explicit condition-based queries, the SFT+RFT model achieved a recall of 0.410, outperforming the SFT-only model (0.391) and the RFT-only model (0.178).
@@ 2.3. Results on User Profile-based Query
Observation 7. For different LLMs, Gemini-1.5 Pro and DeepSeek-R1 demonstrated better performance compared to other models. As shown in Table 5, larger and more advanced models, such as DeepSeek-R1 and Gemini-1.5-Pro, consistently achieve higher Precision and Recall. In contrast, smaller models like GPT-4o-mini tend to lag behind. From the FTR (Failure to Recommend) metrics perspective, most models exhibit a low failure rate, indicating their ability to effectively follow instructions and generate recommendation lists in the required format.
@@
Model Interest-based Query Demographics-based Query Average
@@
Observation 8. For different query types, Demographics-based Query exhibit lower Recall than Interest-based Query but similar Precision. Across all models, Recall for Demographics-based Query is generally lower than for Interest-based Query. This suggests that LLMs struggle to identify relevant recommendations when relying on user demographics. The difference in Recall can be attributed to the distinct ways in which these queries are constructed. Interest-based Query leverages clear interest, making it easier for models to generate relevant recommendations. In contrast, Demographics-based Query requires recommendation assistants to infer preferences from broader demographic and behavioral trends, which introduces more variability and ambiguity.
Observation 9. For Interest-based Query, queries constructed from less prevalent interests tend to be more challenging. As shown in Figure 6, we categorized the queries into four groups based on their popularity and conducted experiments on the movie and book datasets. It can be observed that interests that are shared by more users tend to yield higher Precision and Recall. This is likely because widely shared interests are more easily recognizable and interpretable by LLMs, leading to better performance. In contrast, the book domain shows an opposite trend. This is due to the nature of book data and the evaluation process. Books often have multiple editions or publishers, and popular titles tend to have more variants. Therefore, the popular book (e.g., “Dune”) can suffer lower metrics because LLMs confuse editions/publishers (e.g., “Dune: 1965” vs. “2021 adaptation”). For less popular books, fewer variants exist, making it easier for LLMs to provide exact matches, which leads to higher performance metrics.
@@
Figure 7. Performance of queries with various demographics.
Observation 10. For Demographics-based Query, the performances of LLMs show notable variations based on different user demographics. For instance, as figure 7 shows, we note that models exhibit higher accuracy for female users. It reflects more consistent preference patterns of female users (e.g., stable genre preferences like romantic comedies), which allows the model to make accurate recommendations. From the perspective of age, LLMs perform better regarding Recall for people aged 50–55. This is likely because their preferences are more focused and less affected by rapidly changing popular culture compared to younger users, who tend to have more diverse preferences. For users aged 56 and above, the model’s effectiveness is not as strong as it is for the 50–55 age group, possibly because older users engage less frequently in online activities, such as rating movies. As a result, there is less data about this age group in the training set of LLMs, which makes it more challenging for LLMs to understand their preferences.
Table 6. Fine-tuning on User Profile-based Query
Domain Model Fine-tuning Interest-based Demographics-based
Movie Qwen-2.5-3B - 0.015 0.029
SFT 0.287 0.154
RFT 0.143 0.228
SFT+RFT 0.289 0.230
Llama-3.2-3B - 0.014 0.019
SFT 0.139 0.102
RFT 0.111 0.122
SFT+RFT 0.144 0.131
Book Qwen-2.5-3B - 0.026 -
SFT 0.268 -
RFT 0.108 -
SFT+RFT 0.270 -
Llama-3.2-3B - 0.022 -
SFT 0.189 -
RFT 0.095 -
SFT+RFT 0.193 -
Observation 11. For User Profile-based Query, both fine-tuning methods lead to clear performance gains. As shown in Table 6, all fine-tuned models consistently outperform the base model. The SFT+RFT approach delivers the strongest overall performance, with SFT close behind. However, in certain cases—such as demographics-based queries—RFT outperforms SFT; for example, in the Qwen-2.5-3B model, RFT achieves a recall of 0.228 compared to SFT’s 0.154. We also observe that after fine-tuning, Qwen models generally show better results than the LLaMA models.
@@ 3. Related Work
Recent advances in LLM-based recommender systems (Wang et al., 2025; Qu et al., 2025, 2024) have led to the development of various benchmarks to evaluate their capabilities (Liu et al., 2023; Jiang et al., 2024; Tan et al., 2025; Xue et al., 2025; Zhang et al., 2023). Early efforts like LLMRec (Liu et al., 2023) focus on traditional recommendation tasks (e.g., rating prediction, sequential recommendation), establishing baselines for LLMs against classical methods. In addition to traditional recommendation tasks, recent studies have increasingly focused on evaluating the unique characteristics of LLM for recommendations. For instance, one recent study (Jiang et al., 2024) introduces a multi-dimensional framework that assesses not only utility and novelty but also LLM’s specific behaviors like history length sensitivity (how recommendations perform with sparse user histories) and hallucination (generation of non-existent items). Similarly, PerRecBench (Tan et al., 2025) isolates personalization accuracy by eliminating biases in users and item ratings, testing whether LLMs infer user preferences.
Meanwhile, with the rise of LLMs, varying benchmarks have been proposed to evaluate their capabilities as assistants. For example, Gaia (Mialon et al., 2023) measures LLM’s capabilities as a general-purpose AI assistant. In addition, domain-specific frameworks like VoiceBench (for voice assistant) (Deng et al., 2024) and Mobile-Bench (for mobile assistant) (Chen et al., 2024) assess the ability of LLM as the specific assistant. However, despite progress in both recommendation and assistant evaluation, there remains a gap in assessing LLMs as personalized recommendation assistants. To bridge this gap, we introduce RecBench+, a novel benchmark designed to evaluate LLMs’ potential as personalized recommendation assistants.
@@ 4. Conclusion
In this paper, we introduce a new paradigm for next-generation recommender systems and propose a novel benchmark dataset, RecBench+, to evaluate the potential of LLMs as personalized recommendation assistants. Our benchmark includes five specially designed query types that comprehensively assess LLM-based assistants’ capabilities in handling complex user queries in real-world scenarios. We conduct extensive experiments on a diverse set of LLMs, including those with advanced reasoning capabilities. Through these experiments, we present 11 detailed observations that shed light on the strengths and limitations of LLMs in serving as personalized recommendation assistants. We also explore the impact of different fine-tuning strategies on enhancing recommendation performance. We hope our benchmark and findings lay a foundation for future research and development in this domain.
Acknowledgements.
The research described in this paper has been partially supported by the General Research Funds from the Hong Kong Research Grants Council (project no. PolyU 15207322, 15200023, 15206024, and 15224524), internal research funds from Hong Kong Polytechnic University (project no. P0042693, P0048625, and P0051361). This work was supported by computational resources provided by The Centre for Large AI Models (CLAIM) of The Hong Kong Polytechnic University.
APPENDIX
Appendix A Prompts for Model Testing